Write the equation for the line y - 6 = -6 (x - 6) in the form y = m x + b, and enter your answer in this form.
Answer by Alan3354(69443)

Write the equation for the line y - 6 = -6 (x - 6) in the form y = m x + b
-------------
Solve for y.